OpenStudy (anonymous):

What is the simplified form of:

OpenStudy (anonymous):

\[\sqrt{128n^17}\] To the seventh power .-.

OpenStudy (compassionate):

Is it n^1^7

OpenStudy (anonymous):

n^17

OpenStudy (compassionate):

\[\sqrt{128n^{17}}\]

OpenStudy (anonymous):

Yes. It wouldn't let me write it like that .-.

OpenStudy (compassionate):

Okay, what is the square root of 128?

OpenStudy (compassionate):

g2g

OpenStudy (anonymous):

11.31. and Okay /:

OpenStudy (uri):

Look closely...

OpenStudy (anonymous):

\[8\sqrt{2}\]

OpenStudy (uri):

Ain't you smart?

OpenStudy (anonymous):

Then you input the n^17? which is; \[8n^8\sqrt{2n}\] ?

OpenStudy (uri):

Careful.....

OpenStudy (uri):

Do you have options?

OpenStudy (anonymous):

Yes.

OpenStudy (uri):

Post it.

OpenStudy (anonymous):

\[8n^8\sqrt{2n}\] \[4n^8\sqrt{8n}\] \[2n^8\sqrt{8n}\] \[8n^8\sqrt{4n}\]

OpenStudy (uri):

You got A one right? Go with it.

OpenStudy (anonymous):

Alrighty. Thanks for helping.